### v3

# Multi-Sensor Diffusion-Driven Optical Image Translation for Large-Scale Applications

Abstract
Comparing images captured by disparate sensors is a common challenge in remote sensing. This requires image translation—converting imagery from one sensor domain to another while preserving the original content. Denoising Diffusion Implicit Models (DDIM) are potential state-of-the-art solutions for such domain translation due to their proven superiority in multiple image-to-image translation tasks in classic computer vision. However, these models struggle with large-scale multi-patch imagery, often focusing solely on small patches and resulting in inconsistencies across the full image. To overcome these limitations, we propose a novel method that leverages DDIM for effective optical image translation over large areas. Our approach super-resolves large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. Our contributions lie in new forward and reverse diffusion processes, which are crucial for addressing the challenges of large-scale image translation. Extensive experiments with a dataset of paired Sentinel-II and Planet Dove images demonstrate that our approach provides precise domain adaptation and artifact reduction. The proposed method preserves image content while improving radiometric (color) accuracy and feature representation. The outcome is a high-resolution large-scale image with consistent patches, vital for applications such as heterogeneous change detection (HCD). We present a unique training and testing algorithm rooted in DDIMs, a thorough image quality assessment, and a comparative study against the standard classifier-free guided DDIM framework and five other leading methods. The usefulness of our approach is further demonstrated in HCD tasks performed in the urban settings of Beirut, Lebanon, and Austin, USA.
I Introduction
[…]
Image-to-image translation (I2I) involves transforming an image from one data domain to another while preserving its original content. This technique is useful for applications such as super-resolution, inpainting, image decompression, and domain adaptation [Saharia2021, Dhariwal2021, I2ISeo2023, Park2019, I2IIsmael2023, I2ITasar2020, I2IZhang2023, I2ISokolov2023].
In remote sensing, I2I is valuable when comparing images from different sensors. For example, a deep learning model can translate a low-resolution domain $A$ optical image to a high-resolution domain $B$ optical image.This creates synthetic high-resolution images from low-resolution ones, which are useful when original high-resolution images are unavailable, costly, or affected by weather conditions. However, I2I in this context is challenging due to differences in spatial and spectral characteristics, noise levels, and geometric distortions specific to each sensor. Therefore, an I2I model must preserve the content and structure of the source image while adapting to the radiometric properties and spatial resolution of the target domain.
Despite advancements in image-to-image (I2I) translation for remote sensing, most literature focuses on generating individual synthetic patches for augmenting training data sets used in downstream tasks [I2ITasar2020, I2IIsmael2023, I2ISokolov2023], or on SAR to Optical translation [CDLi2021, I2IZhang2023, I2ISeo2023]. While highly valuable, these methods do not address the challenge of translating large-scale images across disparate optical sensors with different spatial resolutions, particularly ensuring feature consistency between input and output images. This gap is significant as it hinders the fidelity of the translated images, where preserving salient features without introducing non-existent elements is crucial. Additionally, these methods are usually tested on benchmark datasets or small datasets, which may not adequately demonstrate their effectiveness in real-world scenarios.
An appealing choice for addressing these challenges is the use of Denoising Diffusion Models (DDMs) [Ho2020], which have shown exceptional image generation capabilities in other contexts, surpassing adversarial models without common training issues like mode-collapse [mansourifar_i3eaccess2022]. However, standard diffusion methods seem to struggle with high variability in overall patch colorization across samples with similar features in a vast data set. Also, at times they produce inconsistencies in feature generation across neighboring patches due to their stochastic nature [Ho2020].
To overcome these limitations, we propose a novel method that leverages Denoising Diffusion Implicit Models (DDIM) for effective optical image translation over large areas. Our approach super-resolves large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. We address the issues of high variability in patch colorization and inconsistencies in feature generation by designing novel forward and reverse diffusion procedures. These procedures enhance spatial resolution and align radiometric differences between domains, ensuring high-quality, consistent results.
Notably, this work is the first to integrate large-scale optical domain adaptation and super-resolution into a single task using a DDPM-based method. This integration generates large multi-patch images with consistent and color-accurate patches, showing significant visual resolution improvement while preserving the original content (see Section V). The robust training dataset and scalability of our approach contribute to its state-of-the-art performance, making it suitable for practical applications in remote sensing.
We apply our DDM-based method to generate synthetic high-resolution images for eight regions of interest: Beirut, Austin, Tlaquepaque, and five other locations. In our experiments, Sentinel-II imagery ($10\text{m}^{2}/\text{pixel}$) is translated into synthetic versions of Planet Dove imagery ($3\text{m}^{2}/\text{pixel}$). We perform two heterogeneous change detection (HCD) experiments to test the method in a practical application. HCD compares images from different sensors or modes of acquisition to identify changes in a specific area. Valuable for quick disaster damage assessment, achieving accurate and consistent results is challenging due to differences in image characteristics and quality [Chini2023, Touati2020, Zhang2022, WangCD2022, Lv2022, Wu2022].
While we present heterogeneous change detection (HCD) experiments to showcase one application of our image-to-image (I2I) translation algorithm, it is important to note that these experiments are not intended to demonstrate peak performance in HCD. The simplistic image processing algorithm used in these experiments does not exploit the resolution gains from our diffusion-based image translation model. Instead, the purpose is to illustrate the potential use of our I2I translation method in HCD. Developing a learning-based HCD method that fully leverages the added resolution is beyond the scope of this article.
Our contributions are summarized as follows:
  1. 1.
    We propose original training and testing procedures based on denoising diffusion models that translate optical images from a low spatial resolution domain to a higher one over large areas while maintaining high inter-patch and input-output consistency.
  2. 2.
    We present image quality comparisons between our method and other cutting-edge solutions, along with ablation experiments over hundreds of square kilometers of imagery, highlighting the contributions of our approach compared to traditional denoising diffusion.
  3. 3.
    We present heterogeneous change detection tests in Beirut, Lebanon, and Austin, USA, indicating the potential usefulness of our method in HCD tasks.
II Related Works
II-1 Denoising Diffusion Models
Ho et al. [Ho2020] proposed the use of a denoising framework for image generation, the so-called Denoising Diffusion Probabilistic Model (DDPM). It showed incredible image generation quality and overcame most state-of-the-art methods available at the time. However, the inference process was cumbersome since hundreds of model evaluations per image were necessary to reach optimal performance. The work Denoising Diffusion Implicit Models (DDIM) [Song2020] approached this issue by presenting a new interpretation of the reverse diffusion process. Differently from the work in [Ho2020], they treated reverse diffusion as a non-Markovian process by skipping multiple reverse diffusion steps. Furthermore, they experimented with eliminating the randomness of intermediate diffusion steps—which came from the addition of a random noise matrix to each reverse diffusion step result—and noticed that this enabled an even greater reduction of model evaluations during inference at the cost of sample variety. In practice, these combined modifications reduced the necessary number of model evaluations during inference by up to two orders of magnitude while maintaining high-generation quality.
As originally defined in [Ho2020], denoising diffusion models are unconditional, i.e., they are trained to generate images with distributions similar to those it has seen during training without any prior knowledge about how a particular image should resemble or what it should contain. This limited the applicability of such models in multiple condition-guided tasks, such as image inpainting, super-resolution, colorization, and domain adaptation. To make them possible, two works [Nichol2021, Saharia2021] experimented with adding a conditioning matrix as an additional model input. This partially worked since, as observed by Dhariwal et al. [Dhariwal2021], diffusion models have a tendency to ignore the condition and focus only on the noisy input to be denoised. The work in [Dhariwal2021] also proposed a possible strategy to alleviate this issue: the use of a separate classifier through which conditioning information is directly inserted into the reverse diffusion procedure. This proved to be useful, as the images generated with this procedure were strongly linked with the conditions added to the model. Still, the need to have a second classifier model is cumbersome and sometimes impractical. Due to that, Ho et al. [ho2022classifierfree] proposed Classifier-free Guidance, which, as the name implies, removes the need for a separate classifier network that guides the model with conditioning information. They accomplish that by making a small modification in the DDPM training and inference procedures that, in summary, treats the diffusion model as if it consisted of a combination of an unconditional and a conditional model. Hence, during inference, it became possible to determine how much the conditional and unconditional virtual parts of the model influenced the generation and, therefore, how strongly the reverse diffusion should rely on the conditioning information.
II-2 Image-to-Image Translation
In the context of image-to-image translation, Saharia et al. [Saharia2021] investigated the use of DDPMs in different I2I translation tasks: colorization, image restoration, inpainting, and uncropping, obtaining remarkable image generation quality and high sample diversity. Moreover, in the presented super-resolution tests, it beat deep regression-based models by an impressively high margin. The work in [Wang2022] used classifier-free guidance [ho2022classifierfree] in the task of semantic image generation—where the user inputs the information on what each area of the generated image should contain—together with a novel conditional model architecture that includes multi-head attention-based mechanisms. In the considered test sets, they beat all state-of-the-art methods in image quality metrics. Image Super-Resolution via Iterative Refinement [Saharia2021SR3] employed DDPM in the task of super-resolution. They trained a conditional model, referred to as SR3, to generate high-resolution samples starting from their low-resolution versions, beating all compared methods in image quality metrics and fool rate tests.
Other works focused on remote sensing-related tasks. Ismael et al. [I2IIsmael2023] proposed a GAN-based I2I translation method with the goal of performing remote sensing semantic segmentation of optical images from unknown domains using a segmentation network trained with optical images from a single domain. The methods in [I2ITasar2020] and [I2ISokolov2023] proposed style-transferring domain adapting frameworks that generate extra training data matching the appearance of a particular same-resolution domain to alleviate train/test data drift. The method in [CDLi2021] uses a cycle-consistent GAN to translate optical images into SAR images to perform change detection. It showed promising change detection results. However, training and testing data are limited to just a few acquisitions. The method in [I2IZhang2023] is a thermodynamics-inspired translation network that obtained interesting results in SAR-to-Optical translation tasks with its unconventional design. The work in [I2ISeo2023] proposes a DDM-based Brownian Bridge method for SAR-to-Optical I2I translation. There, the authors suggest the use of a modified forward diffusion procedure that, instead of resulting in a pure-noise image, the target SAR image is iteratively transformed into the source optical image. They reason that this could help the model more easily understand the relationships across domains. Xiao et al. propose a diffusion-based super-resolution method that seeks to further explore prior information from the low-resolution input. When tested on a small-scale data set, it performed well.
(a) Beirut Pre Setninel-II
(b) Beirut Pre Planet Dove
(c) Beirut Post Setninel-II
(d) Beirut Post Planet Dove
Figure 1: Surface reflectance imagery from Beirut, Lebanon, of low and high spatial resolutions, captured at similar time intervals, from Sentinel-II (10m) and Planet Dove (3m) sensors. Figures 1a and 1c are the pre- and post-event images from Sentinel-II, respectively. Figures 1b and 1d are the pre- and post-event images from Planet Dove, respectively.
III Proposed Method
Our method addresses the challenging task of large-scale, patch-consistent optical-to-optical image translation from a lower to a higher spatial resolution sensor domain. While spatial resolution is the most noticeable difference, other disparities such as radiometric fidelity, illumination conditions, and sensor spectral sensitivity also increase the radiometric discrepancy between the two domains [CuiColor2020].
To tackle these challenges, we leverage Denoising Probabilistic Diffusion Models (DDPMs) to not only enhance spatial resolution but also align subtle yet significant differences between the two domains, ensuring improved radiometric fidelity and consistency. This requires a training dataset with pairs of co-registered low spatial resolution (LR) domain $A$ and high spatial resolution (HR) domain $B$ images, capturing the same scenes with different sensors for proper domain adaptation.
The method hereby proposed exclusively employs Denoising Diffusion Implicit Models (DDIM) [Song2020], which are advantageous due to their reduced number of model forward passes during inference and their deterministic behavior in intermediate steps. However, our approach differs from the standard classifier-free guided DDIM framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022]. The standard framework struggles with the high variability in patch coloration within large training sets, leading to inaccurate colorization in generated patches.
Additionally, due to the stochastic nature of denoising diffusion models, different initial noise matrices can result in varying outputs for the same input [Ho2020]. This randomness can cause inconsistencies in feature generation across neighboring patches, which is problematic for image-to-image translation in remote sensing.
The proposed method addresses these deficiencies by implementing novel forward and reverse diffusion procedures, ensuring high-quality, consistent results.
The model architecture and the associated hyperparameters used for our experiments are discussed in Appendix A. Details about the denoising diffusion working principles are presented in Appendix B.
Figure 2: Diagram illustrating the training step procedure for the proposed DDM-based image-to-image translation method.
Algorithm 1 Patchwise Training Step
0:  $\bm{I}_{\text{in}}^{\text{LR,local}}$, $\bm{I}_{\text{in}}^{\text{LR,global}}$, $\bm{I}_{\text{out}}^{\text{HR}}$, $\bm{\gamma}$, $p_{\text{uncond}}$, $\bm{\theta}_{\tau}$The model at training step $\tau$}
0:  $\bm{\theta}_{\tau+1}$
1:  $\bm{I}_{\text{in}}^{\text{LR}}=\left[\bm{I}_{\text{in}}^{\text{LR,local}}\>;\bm{I}_{\text{in}}^{\text{LR,global}}\right]$
2:  $u\thicksim\text{Bernoulli}(p_{\text{uncond}})$
3:  if $u\equiv 1$ then
4:     $\bm{x}={\O}_{2n_{\text{ch}}\times h\times w}$
5:  else
6:     $\{\bm{x},\_,\_,\_\}=\mathbf{W}\left(\bm{I}_{\text{in}}^{\text{LR}},2n_{\text{ch}},h,w\right)$
7:  end if
8:  $\{\bm{y}_{0},\_,\_,\_\}=\mathbf{W}\left(\bm{I}_{\text{out}}^{\text{HR}},n_{\text{ch}},h,w\right)$
9:  $\bm{\varepsilon}=\left(\varepsilon_{i,j,k}\right)_{1\leq i\leq n_{\text{ch}},\>1\leq j\leq h,\>1\leq k\leq w}$$\quad\varepsilon_{i,j,k}\sim\mathcal{N}(0,1)\>\forall\>i,j,k$
10:  $\gamma_{t}\sim\text{Uniform}(\bm{\gamma})$
11:  $\bm{y}_{t}\leftarrow\sqrt{\gamma_{t}}\bm{y}_{0}+\sqrt{1-\gamma_{t}}\bm{\varepsilon}$Forward diffusion}
12:  $\hat{\bm{\varepsilon}}\leftarrow\bm{\theta}_{\tau}\left(\bm{y}_{t},\bm{x}\right)$
13:  $\bm{\theta}_{\tau+1}\leftarrow$ Optimization step on $\nabla_{\bm{\theta}_{\tau}}\>\mathcal{L}({\bm{\varepsilon}},\hat{\bm{\varepsilon}})$
Algorithm 2 Image Whitening $\mathbf{W}$
0:  $\bm{I}_{\text{c}}$, $n_{\text{ch}}$, $h$, $w$
0:  $\bm{I}_{\text{w}}$, $\bm{m}_{1},m_{2},m_{3}$
1:  $\bm{I}_{\text{w}}\leftarrow\bm{0}_{n_{\text{ch}}\times w\times h}$
2:  $\bm{m}_{1}\leftarrow\bm{0}_{n_{\text{ch}}}$
3:  $m_{2}\leftarrow 0$
4:  $m_{3}\leftarrow 0$
5:  for $i\leftarrow 0\text{ to }n_{\text{ch}}-1$ do
6:     ${m}_{1}^{i}\leftarrow\mu(\bm{I}_{\text{c}}^{i})$
7:     $\bm{I}_{\text{w}}^{i}\leftarrow\bm{I}_{\text{c}}^{i}-{m}_{1}^{i}\cdot\mathbf{J}_{h\times w}$$\mathbf{J}_{h\times w}$ is the all-ones $h\times w$ matrix}
8:  end for
9:  $m_{2}\leftarrow\min\bm{I}_{\text{w}}$
10:  $\bm{I}_{\text{w}}\leftarrow\bm{I}_{\text{w}}-m_{2}\cdot\mathbf{J}_{n_{\text{ch}}\times h\times w}$
11:  $m_{3}\leftarrow\max\bm{I}_{\text{w}}$
12:  $\bm{I}_{\text{w}}\leftarrow 2\left(\bm{I}_{\text{w}}/m_{3}-0.5\cdot\mathbf{J}_{n_{\text{ch}}\times h\times w}\right)$
III-A Training
Our training procedure diverges from the conventional DDPM by incorporating modifications to the forward diffusion process through Whitening and Coloring techniques. These modifications are critical as they prevent the diffusion network from being confounded by the high variability in patch coloration found within extensive training datasets, particularly when patches have highly similar content. Without these measures, the network fails to accurately predict the channel mean values for the output patches, as shown in Section V. Details of the implementation of DDPM and DDIM are presented in Appendices B-A and B-B.
Algorithm Description
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the similar features. Thus, patches aligned in time are required for training to work. Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$. The matrix $\bm{I}_{\text{in}}^{\text{LR,local}}$ is defined as the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference. $\bm{I}_{\text{in}}^{\text{LR, global}}$ represents the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast patches that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. This is done to give more context to the model of what exactly lies around the input patch. Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix (like in [ho2022classifierfree]). To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until convergence.
Figure 3: Diagram illustrating the inference procedure for the proposed DDM-based image-to-image translation method. Here, the color information variables $\bm{m}_{1}$, $m_{2}$, $m_{3}$ are provided from an external image, e.g., a post-event Planet Dove image.
Algorithm 3 Patchwise Inference
0:  $\bm{\theta},\bm{I}_{\text{in}}^{\text{LR,local}}$, $\bm{I}_{\text{in}}^{\text{LR,global}}$, $n_{\text{ch}}$, $h$, $w$, $\omega_{\text{uncond}}$, $N_{\text{DDIM}}$, $d$, $n_{\text{noisy}}$ Optional: $\bm{m}_{1}$, $m_{2}$, $m_{3}$Color information variables.}
0:  $\bm{I}_{\text{out}}^{\text{HR,local}}$
1:  $\bm{I}_{\text{in}}^{\text{LR}}=\left[\bm{I}_{\text{in}}^{\text{LR,local}}\>\bm{I}_{\text{in}}^{\text{LR,global}}\right]^{T}$
2:  if $\bm{m}_{1}$, $m_{2}$, $m_{3}$ are all provided then
3:     $\{\bm{x}_{\text{local}},\_,\_,\_\}=\mathbf{W}\left(\bm{I}_{\text{in}}^{\text{LR,local}},n_{\text{ch}},w,h\right)$
4:     $\{\bm{x}_{\text{global}},\_,\_,\_\}=\mathbf{W}\left(\bm{I}_{\text{in}}^{\text{LR,global}},n_{\text{ch}},w,h\right)$
5:  else
6:     $\{\bm{x}_{\text{local}},\bm{m}_{1},m_{2},m_{3}\}=\mathbf{W}\left(\bm{I}_{\text{in}}^{\text{LR,local}},n_{\text{ch}},w,h\right)$
7:     $\{\bm{x}_{\text{global}},\_,\_,\_\}=\mathbf{W}\left(\bm{I}_{\text{in}}^{\text{LR,global}},n_{\text{ch}},w,h\right)$
8:  end if
9:  $\bm{x}=\left[\bm{x}_{\text{local}};\>\bm{x}_{\text{global}}\right]$
[…]
11:  $\hat{\bm{Y}}_{0,\text{lowqual}}\leftarrow\mathbf{0}_{n_{\text{noisy}}\times n_{\text{ch}}\times h\times w}$
12:  for $i\leftarrow 1\text{ to }n_{\text{noisy}}$ do
[…]
14:  end for
[…]
17:  $\bm{I}_{\text{out}}^{\text{HR,local}}=\mathbf{C}(\hat{\bm{y}}_{0},n_{\text{ch}},h,w,\bm{m}_{1},m_{2},m_{3})$
III-B Inference
The inference procedure illustrated by Figure 3 aims to maximize the image generation quality by selecting the best possible initial noise matrix to be fed to the reverse diffusion process. We propose a novel reverse diffusion procedure by integrating the same Whitening and Coloring techniques used during training. To further enhance the fidelity of the generated patches to their inputs, we refine the selection process for the initial noise condition input for the reverse diffusion process, ensuring optimal consistency and translation quality. Implementation details for the DDIM inference process are presented in Appendix B-B.
In our experiments, we observed that, for the same conditioning input $\bm{x}$, the quality of DDIM reverse diffusion outputs generated with different initial noise matrices varied significantly. That is, for a combination of independent and identically distributed random initial noise matrices, fixing $\bm{x}$, the DDIM, in some cases, generates samples entirely consistent with $\bm{x}$, whereas sometimes it generates inconsistent samples with undesired hallucinations. Since the DDIM reverse diffusion is a deterministic process, i.e., given fixed $\bm{\varepsilon}_{T}$ and $\bm{x}$, its output remains the same, the cause for varying consistency levels lies only in the different sampled initial noise matrices, which are the source of the randomness of the whole inference process. Thus, we propose a procedure that seeks to select an initial noise matrix such that the DDIM output possesses the highest consistency compared to $\bm{x}$. This procedure consists of repeating the DDIM reverse diffusion process with a reduced number of iterations, using as input $n_{\text{noisy}}$ different initial noise matrices. The resulting outputs are tested for consistency by individually calculating the peak signal-to-noise ratio (PSNR) between them and $\bm{x}_{\text{local}}$. The one that was responsible for the maximum PSNR is used as the definitive starting noise matrix for the inference process, which takes place with a higher number of iterations to ensure maximal generation quality.
Algorithm Description
[…]
0:  $\bm{I}_{\text{w}}$, $n_{\text{ch}}$, $h$, $w$, $\bm{m}_{1}$, $m_{2}$, $m_{3}$
0:  $\bm{I}_{\text{c}}$
1:  $\bm{I}_{\text{c}}\leftarrow\bm{0}_{n_{\text{ch}}\times h\times w}$
2:  $\bm{I}_{\text{w}}\leftarrow\frac{\bm{I}_{\text{w}}-\min\bm{I}_{\text{w}}\mathbf{J}_{n_{\text{ch}}\times h\times w}}{\max\bm{I}_{\text{w}}-\min\bm{I}_{\text{w}}}$
3:  $\bm{I}_{\text{w}}\leftarrow m_{3}\cdot\bm{I}_{\text{w}}$
4:  $\bm{I}_{\text{w}}\leftarrow\bm{I}_{\text{w}}+m_{2}\cdot\mathbf{J}_{n_{\text{ch}}\times h\times w}$
5:  for $i\leftarrow 0\text{ to }n_{\text{ch}}-1$ do
6:     $\bm{I}_{\text{c}}^{i}\leftarrow\bm{I}_{\text{w}}^{i}+m_{1}^{i}\cdot\mathbf{J}_{h\times w}$
7:  end for
Algorithm 4 Image Colorization $\mathbf{C}$
IV Dataset
For the proposed experiments, paired images from satellites Sentinel-II and Planet Dove were gathered. The former is a widely known open-access optical satellite with a spatial resolution of $10\text{m}^{2}/\text{pixel}$, whereas the latter is a class of commercial satellites aimed at producing high-quality images at a spatial resolution of $3\text{m}^{2}/\text{pixel}$. A dataset containing 77 pairs of surface reflectance rasters has been assembled for training. These rasters are parts of or a combination of acquisitions that were captured at similar time periods. They are properly geolocated, georeferenced, and checked for image quality issues. In total, a vast area of 7894 $7894\text{ km}^{2}$ is covered by the training set, which is comparable to the area of Puerto Rico (9104 $9104\text{ km}^{2}$). As the test set for the image-to-image translation algorithm, 8 pairs of images were selected, which, in total, cover an area of 707 $707\text{ km}^{2}$. With these, image quality metrics are extracted for the proposed method and compared to other possible approaches for the I2I translation task.
For evaluation of the performance of our method in a change detection task, we use pre and post-images from the port region in Beirut, Lebanon. On August 4th, 2020, an explosion dramatically affected the port and surrounding areas. As they have recovered over time, many changes took place, e.g., new and repaired constructions and structures. Thus, for the change detection test, we use as pre-event images a pair of images acquired hours after the explosion, between 04 and 20 of August 2020, when the effects of the event are still evident. As post images, acquisitions from July 2023 of the exact same region are used. Figure 1 shows the Beirut pre and post-event images from Sentinel-II and Planet Dove. We also perform change detection tests in a region in Austin, USA, where many urban changes can be observed between the selected pre- and post-event images. They were captured in July 2021 and 2023, respectively.
(a) Sentinel-II
(b) Planet Dove (Post)
(c) Proposed: DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting
(d) DDIM + $\mathbf{W},\mathbf{C}$
(e) DDIM
(f) Regression
(g) Regression + $\mathbf{W},\mathbf{C}$
(h) Pix2Pix [Pix2Pix2017]
(i) ShuffleMixer [sun2022shufflemixer]
(j) SRDenseNet [zhang2018residual]
(k) SwinIR [liang2021swinir]
(l) SR3 [Saharia2021SR3]
Figure 4: Comparison among images generated by the tested I2I methods (c)-(l) and the original Sentinel-II (a) and Planet Dove (b) images. The image generated with the proposed model (c) displays higher resolution features not observable among the others, while maintaining patch-wise feature consistency and avoiding blurriness.
V Experiments
V-A Image Quality Metrics and Ablation Experiments
TABLE I: Image Quality Metrics Comparison with the State-of-the-Art Using $N_{\text{final}}=64$, $N_{\text{pre}}=8$, $n_{\text{noisy}}=8$
[…]
Sentinel-II (Input Images) 0.2977 98.08 12.76
Regression 1 0.2797 64.12 16.72
Regression + $\mathbf{W},\mathbf{C}$ 1 0.2491 67.04 17.48
Conditional DDIM 1 0.4769 87.93 9.919
Conditional DDIM + $\mathbf{W},\mathbf{C}$ 1 0.1993 46.89 15.39
Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting (Proposed) 0.1884 45.64 15.72
Other State of the Art Models
Pix2Pix [Pix2Pix2017] 0.3211 105.0 14.87
ShuffleMixer [sun2022shufflemixer] 0.2249 73.82 16.06
SRDenseNet [zhang2018residual] 0.2296 57.70 17.12
SwinIR [liang2021swinir] 0.2325 56.36 16.68
SR3 [Saharia2021SR3] 0.6235 152.6 9.906
1 Identical neural network architecture as the proposed method.
TABLE II: Comparison Metrics for Different Inference Hyperparameters
[…]
16 2 2 20 0.2229 52.46 14.60
32 4 4 48 0.1967 45.17 15.29
48 6 6 84 0.2013 46.49 15.15
64 8 8 128 0.1884 45.64 15.72
Figure 5: Figures of the region of interest in Tlaquepaque, Mexico. In the first row, images of a far view of the city are displayed. In the second row, we can observe a cropped part of the images in the first row for closer inspection. At the third row, a cropped part of the images from the second row is displayed, for the inspection of meter-scale features. The first, second, and third columns contain images from Planet Dove, Sentinel-II, and Synthetic Planet domains, respectively.
To effectively demonstrate the performance and advantages of the proposed I2I algorithm, a comparison in terms of image quality metrics is proposed. Table I displays the values of three image quality/comparison metrics: Frechét Inception Distance (FID), mean Peak Signal-to-Noise Ratio (PSNR), and mean Learned Perceptual Image Patch Similarity (LPIPS). To obtain these values, we use the test set described in Section IV. The FID [Heusel2017] measures the similarity between two data sets of images—usually a data set of generated images and one of reference images—by computing the distance between their feature representations extracted by a pre-trained Inception network. A lower FID indicates better feature alignment of the generated images with the reference data set. The PSNR [Hore2010] measures the reconstruction error between two images in terms of the ratio of the maximum possible pixel value and the mean squared error. A higher PSNR indicates a lower pixel-wise reconstruction error. It is important to note that PSNR is not a reliable metric for perceptual quality, as it does not account for the human visual system and the semantic content of the images [Saharia2021]. Nonetheless, in the presented results, we compute the mean PSNR (mPSNR) of all Planet and Synthetic Planet corresponding patches, to further enrich the discussion. The LPIPS [Zhang2018LPIPS] measures the perceptual similarity between a pair of images by computing the distance between their feature representations extracted by a pre-trained deep neural network. A lower LPIPS indicates higher perceptual similarity and better visual quality. LPIPS is aligned with human perception, as it captures the semantic and structural differences between images [Zhang2018LPIPS]. As with the PSNR, we compute the mean LPIPS (mLPIPS) of all corresponding Planet and Synthetic Planet patches. In a patch-wise comparison, the mLPIPS metric is better suited as a comparison metric than FID since its calculation is done on a patch-by-patch basis instead of focusing on the whole test set at once. Thus, since our tests involve direct comparison between corresponding images, mLPIPS is also employed.
We compare our proposed I2I algorithm with other possible I2I solutions. First, the Sentinel-II test set is directly compared to Planet Dove images without any I2I translation method. The obtained metrics are shown in the first row of Table I. Second, we train two deep regression models [Lathuiliere2020] whose loss objective is to directly match the pixel values of Planet patches when Sentinel-II corresponding patches are input. They share the same neural network backbone as the proposed DDM-based method. In other words, these models are identical to the one used in the proposed method, with the exception that the output linear layer directly predicts the values of pixels of the corresponding Planet patch for a given input. The first deep regression model, whose metrics are displayed in the second row of Table I, is a conventional regression model, i.e., it does not apply any pre or post-processing to its inputs and outputs apart from simple $[-1,1]$ normalization. The second regression model, whose results are displayed in the third row, is identical to the previous model, except that the whitening and coloring algorithms, defined in Algorithm 2 and 4, are applied to its inputs and outputs, respectively. In the table, we refer to this as Regression + $\mathbf{W},\mathbf{C}$. The goal of testing this configuration is to see what would be the influence of these processing methods in a conventional regression model and whether or not it has a significant impact on the metrics. Following, we test the standard classifier-free guided DDIM [ho2022classifierfree], i.e., without the proposed whitening, coloring, and PSNR voting processes, using the exact same model architecture as the proposed method, to perform an ablation of these techniques. It is referred in Table I as Conditional DDIM. Then, we train an identical model, now adding the whitening and coloring procedures. To measure the impact of the PSNR voting in the inference results, the image quality metrics for this model are separated into two rows, the first being without PSNR voting—denoted as Conditional DDIM + $\mathbf{W},\mathbf{C}$—and the second with it—Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. All DDIM results were obtained using $N_{\text{final}}=64$. The proposed method, which includes PSNR voting, used $N_{\text{pre}}=8$ and $n_{\text{noisy}}=8$.
We also compare the performance of our method with five other state-of-the-art approaches for super resolution and image-to-image translation tasks. The first method is the well-established Pix2Pix [Pix2Pix2017], which has been proposed for multiple generative tasks, including image-to-image translation. It is an adversarial framework that uses both regression and discriminative losses to guide training of a generator UNet-like network. Following, we compare our method against the SwinIR [liang2021swinir] super resolution model, which leverages the Swin Transformer architecture for enhanced image processing capabilities. The ShuffleMixer [sun2022shufflemixer] architecture, known for its efficient mixing of features through channel shuffling and spatial mixing, is another model we compare against. The SRDenseNet [zhang2018residual] is also included in our comparison—a model that utilizes dense connections to improve the flow of information and gradients throughout the network, which is particularly beneficial for the task of super resolution. Finally, we present results for the SR3 super resolution diffusion-based framework [Saharia2021SR3], using DDIM sampling with 64 steps. The UNet structure of SR3 is analogous to the one used in the article that introduced the DDPM procedure [Ho2020]. Thus, comparing it to our method provides an insight on the advantages of the neural network structure chosen for our experiments.
V-A1 Comparison Metrics
Looking at the numbers of Table I, we observe that the full proposed technique, Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting, reaches, by a good margin, better results in mLPIPS and FID metrics when compared to all other methods and to the Sentinel-II baseline. FID dropped by 52.44 points when compared to Sentinel-II images—an expressive improvement of $53.46\%$. When compared to Regression and Regression + $\mathbf{W}$, the FID has been respectively reduced by 18.48 and by 21.8 points—an improvement of $28.82\%$ and $31.92\%$. Among the state-of-the-art methods, only SwinIR and SRDenseNet reached comparable mLIPS and FID values to the ones from our approach.
Interestingly, the vanilla DDIM-based models (Conditional DDIM and SR3) lose to all methods in any of the metrics. Even worse, they lose to the original images in both mLPIPS and mPSNR. The empirical reason for this is that the vanilla diffusion training seems to get confused by the multiple different overall tonalities of patches containing similar features, which impacts the convergence ability of the model. Deck and Bischoff [deck2023easing] noticed that diffusion models tend to present such color shifting problems, which tends to get worse for larger images. This issue seems to be eliminated when whitening and coloring procedures are used, as evidenced by the performance superiority of both Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. Looking at mLPIPS results, we observe that the regression-based models beat by a small margin the original Sentinel-II images, whereas Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting show a much more significant improvement in the same regard. This indicates that, in a patch-by-patch basis, only DDIM + $\mathbf{W},\mathbf{C}$ and DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting models were able to heavily reduce the perceptual distance of patches when compared to their Planet Dove versions. The proposed method with PSNR Voting has reduced mLPIPS by 0.1093 points, meaning that the produced patches are, on average, $36.71\%$ more perceptually similar to their corresponding Planet Dove patches than their Sentinel-II versions. When we compare the last two rows of the table, PSNR voting brings a significant improvement in all metrics: $-5.469\%$ for mLPIPS, $-2.665\%$ for FID, and $+2.144\%$ for mPSNR. We observed, moreover, that PSNR voting is specially useful to discard rare but existent undesired model hallucinations and poor translations, which would negatively impact, for example, change detection results. These improvements, due to the increased number of model evaluations, come at the cost of higher numerical complexity.
While regression models achieve higher mPSNR values, as discussed, PSNR’s preference for blurry regression-produced outputs does not align well with human perception. The same can be said about the SwinIR, ShuffleMixer and SRDenseNet methods, as they are also trained using regression objectives. Consequently, learning-based metrics, like the FID and the LPIPS, lately have been the standard in image generation comparisons, rendering the PSNR performance superiority of regression models irrelevant.
As discussed, the very poor performance shown by the SR3 model is in part due to what plagued the performance of the classically trained DDIM model. However, even compared to it, the SR3 model presented far worse numbers. This fact shows how the model architecture chosen for our method has been able to better extract the complex features of the conditioning input and generate better translations.
V-A2 Inference Hyperparameters
To verify the impact of the choice of inference hyperparameters linked to the proposed method, we present a performance comparison for different values of $N_{\text{final}}$, $N_{\text{pre}}$ and $n_{\text{noisy}}$ in Table II. In terms of mLPIPS, $N_{\text{final}}=64$, $N_{\text{pre}}=8$ and $n_{\text{noisy}}=8$ reached the best result of $0.1884$. We use this configuration in our change detection experiments, as a lower perceptual distance between identical patches is crucial for a high change detection performance. Interestingly, however, $N_{\text{final}}=32$, $N_{\text{pre}}=4$ and $n_{\text{noisy}}=4$ reached a lower FID, which suggests that further increasing the number of iterations does not bring any value with regards to image quality. Moreover, unintuitively, $N_{\text{final}}=48$, $N_{\text{pre}}=6$ and $n_{\text{noisy}}=6$ showed worse performance metrics compared to the former configuration. This may be due to the fact that the number of training diffusion steps was set to 1024, and 48 is not a divisor of such number. This implicates that the particular timestep values input to the model during the DDIM sampling were not seen during training, leading to such worse performance values. Finally, for $N_{\text{final}}=16$, $N_{\text{pre}}=2$ and $n_{\text{noisy}}=2$, the performance metrics were clearly degraded, which suggests that $N_{\text{final}}=32$, $N_{\text{pre}}=4$ and $n_{\text{noisy}}=4$ would be a good balance between numerical complexity and performance. In addition, the usefulness of PSNR voting is confirmed when this configuration is compared to the results of the model using 64 DDIM iterations without PSNR voting, displayed in Table I. That is, even with a lower total number of iterations (48 vs 64), the configuration $N_{\text{final}}=32$, $N_{\text{pre}}=4$ and $n_{\text{noisy}}=4$ reached better mLPIPS and FID values compared to the DDIM without PSNR voting.
(a) Pre-Event Sentinel-II
(b) Pre-Event Synthetic Planet Dove (Regression + $\mathbf{W},\mathbf{C}$)
(c) Pre-Event Synthetic Planet Dove (Proposed)
(d) Pre-Event Planet Dove
(e) Post-Event Planet Dove
(f) Change Detection Ground Truth
Figure 6: Figures of the region of interest in the port of Beirut.
V-B Visual Examples
In our evaluation of image generation methods, as detailed in Table I, Figure 4 showcases the visual differences using examples from a region in Beirut. The sequence begins with our proposed method (4c), which utilizes the DDIM approach enhanced by whitening and coloring processes ($\mathbf{W},\mathbf{C}$) and PSNR voting, and ends with the SR3 diffusion-based method (4l). The Sentinel-II (4a) and Planet Dove (4b) images serve as input and target references, respectively.
Our method (4c) demonstrates superior patch consistency and detail resolution, outperforming the other techniques. The addition of $\mathbf{W},\mathbf{C}$ significantly enhances contrast and clarity, as evidenced by the comparison with the vanilla DDIM result (4e), which suffers from tonal inconsistencies and a mosaic-like appearance. These issues are notably absent in Figure 4d, where $\mathbf{W},\mathbf{C}$ have been applied.
Despite these improvements, some regions in Figure 4d exhibit deficiencies. The center-west patch is notably blurry and washed-out, suggesting a failure in the generation process. The south-west and south-east patches display a darker tone compared to adjacent areas, with the south-east patch also presenting blurry, low-resolution features not observed in the corresponding region of Figure 4c. This demonstrates how the PSNR voting mechanism improves overall generation quality.
The regression-based methods (4f and 4g), which share the same neural network backbone as the proposed method, yield images with a blurry appearance, which does not significantly enhance the apparent resolution when compared to the Sentinel-II image.
In a similar venue, the algorithms Pix2Pix (4h), ShuffleMixer (4i), SRDenseNet (4j), and SwinIR (4k) maintain good inter-patch consistency but fall short in capturing the high-resolution details that our method achieves. The SR3 method (4l) performs poorly, with color inconsistencies and feature representation problems that are more severe than those observed when our method is applied without $\mathbf{W},\mathbf{C}$.
(a) Pre-Event Sentinel-II
(b) Pre-Event Synthetic Planet Dove (Regression + $\mathbf{W,C}$)
(c) Pre-Event Synthetic Planet Dove (Proposed)
(d) Pre-Event Planet Dove
(e) Post-Event Planet Dove
(f) Change Detection Ground Truth
Figure 7: Figures of the region of interest in Austin.
In a direct comparison between Sentinel-II, Planet Dove, and the synthetic Planet Dove image generated by the proposed method, Figure 5 presents a region from the city of Tlaquepaque, Mexico. Each of the three rows of figures shows a different view of the city, with different scales, to better compare different aspects of the images. The three columns show images from Planet Dove, Sentinel-II, and images generated by the proposed method, respectively. The first row shows a far view of the city that depicts multiple complex urban features: houses, buildings, storage facilities, factories, roads, and scattered vegetation. When looking at the quality of the figures, Figure 5b clearly shows great contrast between features due to it coming from a higher spatial resolution satellite when compared to Figure 5a. Figure 5c is the synthetic high-resolution image generated with the proposed model using Figure 5a as input. It exhibits a higher contrast between features, compared to the Sentinel-II version, while maintaining feature consistency between patches to the point that it becomes almost impossible to distinguish individual patches in the full images. The same can be said about Figure 5f, which shows a close-up view of a region located around the center of Figure 5c, where homogeneity is still observed, and the super-resolution aspect of the method starts to be noticeable. That is, high-frequency information not clearly observable in Figure 5d gets enhanced by the algorithm, resulting in Figure 5f. Finally, the last row depicts a highly zoomed-up area of the city, where individual residential buildings can be visualized. It is clearly noticeable in Figure 5h how the higher spatial resolution information provided by the Planet Dove sensor enables the visualization of details otherwise imperceptible in the Sentinel-II image (Figure 5g), as well as properly separates features coming from different elements in the image, e.g., roads are easily distinguishable from buildings, whereas in 5g the pixels from narrow roads are interleaved with pixels from buildings. This drawback of Sentinel-II images is alleviated by the proposed method, as can be observed in Figure 5i. There, pixel-level details are almost as clear and as distinguishable as in Figure 5h. In addition to this, the synthetic image does not show visible undesirable hallucinations, which enables change detection tasks.
Figures 6 and 7 respectively show images from the selected regions of Beirut and Austin. Their first, second, third, and fourth rows individually show pre-event Sentinel-II, pre-event Synthetic Planet Dove (Regression + $\mathbf{W}$), pre-event Synthetic Planet Dove (Proposed), pre-event Planet Dove, post-event Planet Dove, and the change detection ground truth obtained by applying Algorithm 5, which is discussed in Section V-C, using Planet Dove pre and post-event images. Instead of manual annotations, we use this ground truth since our I2I translation method aims to closely match low-resolution input to the style and resolution of high-resolution images, specifically Planet Dove in our experiments.
In both regions, the tonality of the synthetic images generated using the proposed method (Figs. 6c and 7c) closely resembles those of the Planet pre-event images (Figs. 6d and 7d). Also, when compared with the pre-event Sentinel-II images used as inputs to the DDM-based method (Figs. 6a and 7a), it is easily observable how they have an improved apparent resolution and contrast. It is also noticeable how no perceptible undesired hallucinated features are present, favoring their use for change detection. Compared to the Regression + $\mathbf{W}$ synthetic images (Figs. 6b and 7b), our method delivered images with higher contrast and visual resolution, which translated into images that better match the characteristics of Planet Dove imagery.
The few patching brightness mismatches can be easily alleviated by performing inference with overlapping patches, as done in [wan_psc_2024]. However, this would greatly increase the numerical complexity of the inference, e.g. by a factor of 4 if the overlap is 1/2 of the size of the patch, or by a factor of 16 if the overlap is 3/4 of the size of the patch. The reverse diffusion process is costly because it needs numerous model forward passes to infer a single patch. Therefore, we do not explore this possibility in our paper.
Figure 8: Visualization of the intermediate steps of the diffusion process. The first row shows the progression of the noisy image $\bm{y}_{t}$ along diffusion timesteps. The second row shows their corresponding denoised predictions $\bm{y}_{0,t}$.
To demonstrate the dynamics of the forward and reverse diffusion processes for this specific task, Figure 8 shows an example of the intermediate steps of the denoising procedure. The first row shows the progression of the noisy image $\bm{y}_{t}$ along diffusion steps $t\in{0,1,...,T-1}$, whereas the second row shows their corresponding denoised outputs $\bm{y}_{0,t}$. During the very initial steps, most of the low-frequency information is recovered, resulting in blurry low-quality outputs. Around halfway of the process, the denoised image contains almost all high-frequency features of the final prediction. Thus, subsequent steps serve only to recover the finest details for the final prediction.
Algorithm 5 Targetless Change Detection
0:  $\bm{I}_{\text{pre}}$, $\bm{I}_{\text{post}}$, $l$, $n_{\text{ch}}$, $h$, $w$, $\omega$, $w_{\text{gauss}}$, $w_{\text{otsu}},e_{\text{max}},n_{\text{min}}$
0:  $\bm{I}_{\text{diff}}$
1:  $\bm{I}_{\text{pre}}\leftarrow$ $\left[\min\left(I_{\text{pre}}^{i,j,k},\mu\left(\bm{I}_{\text{pre}}\right)+6\sigma_{\bm{I}_{\text{pre}}}\right)\right]_{1\leq i\leq n_{\text{ch}},\>,1\leq j\leq h,\>1\leq k\leq w}$
2:  $\bm{I}_{\text{post}}\leftarrow$ $\left[\min\left(I_{\text{post}}^{i,j,k},\mu\left(\bm{I}_{\text{post}}\right)+6\sigma_{\bm{I}_{\text{post}}}\right)\right]_{1\leq i\leq n_{\text{ch}},\>,1\leq j\leq h,\>1\leq k\leq w}$
3:  $\bm{I}_{\text{pre}}\leftarrow\text{GaussianBlur}\left(\frac{\bm{I}_{\text{pre}}-\min\bm{I}_{\text{pre}}\cdot\mathbf{J}_{n_{\text{ch}}\times h\times w}}{\max\bm{I}_{\text{pre}}-\min\bm{I}_{\text{pre}}},w_{\text{gauss}}\right)$
4:  $\bm{I}_{\text{post}}\leftarrow\text{GaussianBlur}\left(\frac{\bm{I}_{\text{post}}-\min\bm{I}_{\text{post}}\cdot\mathbf{J}_{n_{\text{ch}}\times h\times w}}{\max\bm{I}_{\text{post}}-\min\bm{I}_{\text{post}}},w_{\text{gauss}}\right)$
5:  $\bm{I}_{\text{pre}}\leftarrow\frac{\bm{I}_{\text{pre}}-\mu\left(\bm{I}_{\text{pre}}\right)\mathbf{J}_{n_{\text{ch}}\times h\times w}}{\sigma\left(\bm{I}_{\text{pre}}\right)}$
6:  $\bm{I}_{\text{post}}\leftarrow\frac{\bm{I}_{\text{post}}-\mu\left(\bm{I}_{\text{post}}\right)\mathbf{J}_{n_{\text{ch}}\times h\times w}}{\sigma\left(\bm{I}_{\text{post}}\right)}$
7:  $\bm{I}_{\text{diff}}\leftarrow\left(\bm{I}_{\text{post}}-\bm{I}_{\text{pre}}\right)^{\odot 2}$
8:  $\bm{I}_{\text{diff}}\leftarrow\frac{1}{n_{\text{ch}}}\sum_{i=1}^{n_{\text{ch}}}\bm{I}^{i}_{\text{diff}}$
9:  $\bm{I}_{\text{diff}}\leftarrow\frac{\bm{I}_{\text{diff}}-\min\bm{I}_{\text{diff}}\cdot\mathbf{J}_{n_{\text{ch}}\times h\times w}}{\max\bm{I}_{\text{diff}}-\min\bm{I}_{\text{diff}}}$
10:  $\bm{I}_{\text{diff}}\leftarrow\mathbf{H}\left({\bm{I}}_{\text{diff}}-\omega\mathbf{J}_{n_{\text{ch}}\times h\times w}\right)\odot\bm{I}_{\text{diff}}$
11:  $\bm{I}_{\text{diff}}\leftarrow\text{Otsu}\left(\bm{I}_{\text{diff}},w_{\text{otsu}}\right)$
12:  $\bm{L}_{\text{diff}}\leftarrow\text{DBSCAN}\left(\bm{I}_{\text{diff}},e_{\text{max}},n_{\text{min}}\right)$
13:  $\bm{I}_{\text{diff}}\leftarrow\begin{cases}I_{\text{diff}}^{i,j,k}&\text{if }L_{\text{diff}}^{i,j,k}\neq-1\\ 0&\text{otherwise}\end{cases}$    ${1\leq i\leq n_{\text{ch}},\>,1\leq j\leq h,\>1\leq k\leq w}$
V-C Change Detection as a Use Case
By transforming images from one domain to another, image-to-image translation methods can effectively highlight differences and changes between images taken at different times or under different conditions. Here, we explore the use of our image-to-image translation method for change detection, demonstrating its practical utility in this specific context. It is important to note that change detection is just one possible use case for our method. The results presented here aim to showcase how our image-to-image translation approach can be beneficial for practical applications, without limiting its potential to other domains.
V-C1 Rationale for Selective Comparisons
We focus on evaluating our model’s performance against one of the regression models previously compared: Regression + $\mathbf{W,C}$. Previous experiments have already demonstrated the superiority of our model in image generation compared to state-of-the-art algorithms. Specifically, our method outperforms others in image quality, particularly with respect to the LPIPS metric, which is a patch-by-patch comparative metric. A lower LPIPS score indicates better translation and fewer artifacts, which correlates with higher quality change detection results. To maintain a concise and focused manuscript, we limit our change detection experiments to comparisons against regression models. This approach allows us to highlight the power of using diffusion models against a regression-based objective, which is also employed by the methods ShuffleMixer, SRDenseNet, and SwinIR.
It is important to note that the tested change detection algorithm does not necessarily extract the full potential of our image translation method, as it contains steps that do not fully utilize the added visual resolution to our advantage. Additionally, using this specific simple HCD method with images generated by the proposed method does not necessarily produce better HCD results when compared to other methods such as ShuffleMixer, SRDenseNet, and SwinIR. These methods have performed poorly in super resolution but very well in patch-wise feature consistency, and Algorithm 5 does not make use of the higher resolution features produced by our image translation method. However, employing a more advanced change detection algorithm is beyond the scope of this paper. The HCD experiments are intended to demonstrate that our proposed method could be potentially useful for change detection. A more advanced learning-based change detection procedure using the produced synthetic images will be explored in future work. Furthermore, Regression + $\mathbf{W,C}$ is trained with the same neural network backbone as the proposed method. By doing so, we effectively showcase the advantages of our diffusion-based method under consistent conditions, ensuring a clear and concise presentation of our results.
V-C2 Procedure
Algorithm 5 defines a simple change detection procedure for a pair of images, $\textbf{I}_{\text{pre}}$ and $\textbf{I}_{\text{post}}$, taken before and after a change event, respectively. It is targetless, meaning that no specific class of detections is to be highlighted, but instead, all changes are treated equally. The algorithm, which is used in the change detection experiments presented henceforth, is a structured approach to targetless change detection, leveraging the robustness of Otsu’s method [Crapsu1979], a well-known adaptive thresholding technique. It works by automatically selecting the optimal threshold value to separate pixels into two classes, foreground, and background, based on their grayscale levels, in a sliding-window fashion. Intuitively, it calculates the threshold that minimizes intra-class variance, effectively distinguishing between changed and unchanged areas in an image.
Lines 1 and 2 of Algorithm 5 individually clip $\textbf{I}_{\text{pre}}$ and $\textbf{I}_{\text{post}}$ to their mean plus six times their standard deviations, to remove extreme outliers that could affect the difference image calculation. Lines 3 and 4 apply Gaussian filtering to $[0,1]$-normalized versions of the images, to smooth the images, which alleviates the value transitions, suppressing potential false alarms. Following, in Lines 5 and 6, the images are standardized so that the pixel values of both images become comparable. Line 7 computes the squared difference between the two images in order to highlight the changes, followed by a channel-wise mean in Line 8, to compact the change map into a 2D matrix. Line 10 zeroes out elements of the difference image that are smaller than $\omega$. $\mathbf{H}$ is the element-wise Heaviside step function. By setting such a global threshold, we effectively filter out minor variations in pixel values that are likely due to noise. This pre-processing step simplifies the image data, ensuring that when Otsu’s method is applied, it can more accurately focus on the substantial changes—those that are of real interest in the context of the event being analyzed. Then, in Line 11, the Otsu’s adaptive thresholding [Crapsu1979] is applied with window size of $w_{\text{otsu}}$, to generate a binary map whose nonzero elements correspond to potential changes. In Lines 12 and 13, to alleviate the presence of detection noise, i.e., to remove positive pixels that are completely isolated and scattered through the binary map, Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [DBSCAN1996] algorithm is applied. It locates clusters of positive pixels of significant size and discards pixels that do not belong to any detected cluster, classifying them hereby as noisy pixels. Its parameters are the maximum Euclidean distance between pixels of a cluster, $e_{\text{max}}$, and the minimum number of pixels that a cluster should contain, $n_{\text{min}}$. The output of Line 12, $\bm{L}_{\text{diff}}$, is a matrix with the dimensions of $\bm{\textit{I}}_{\text{diff}}$, whose pixels are integers from -1 to the number of identified clusters. All pixels equal to -1 are classified as noise and are finally discarded in the last line of Algorithm 5, resulting in a clean binary map, where nonzero pixels are classified as changed pixels. For the executed experiments, the values of $w_{\text{gauss}}$, $w_{\text{otsu}}$, $e_{\text{max}}$, and $n_{\text{min}}$ were set to 11, 1023, 5, and 48, respectively.
(a) Sentinel-II (Pre) + Planet (Post)
(b) Regression + $\mathbf{W,C}$ Synthetic Planet (Pre) + Planet (Post)
(c) Proposed Synthetic Planet (Pre) + Planet (Post)
(d) Ground Truth—Planet Pre and Post
Figure 9: Change detection results in Beirut Port Region for a fixed detection rate of 0.75. Pixels in green, blue, and red are true positives, false negatives, and false positives, respectively.
V-C3 HCD Experiments
Two pairs of Sentinel-II pre-event and Planet Dove post-event images from two different cities were selected for a change detection comparison experiment. We also use their respective pre-event Planet Dove images to produce a change detection ground truth. The Sentinel-II pre-event images are part of the test set mentioned in Section IV. The first pair of images is in Beirut, Lebanon, surrounding the port region. It has been extracted from the northern part of the images in Figure 1, where no clouds are present. This image has been considered because it showcases multiple different types of changes, making targetless change detection useful. Moreover, it is a good example to test the robustness of the proposed method. The second pair of images is in Austin, USA, where many construction and terrain-related changes are present. The metrics used henceforth are the detection rate (DR) (recall) and the false alarm rate (FAR), which is defined as the ratio between the number of pixels misclassified as changes and the total number of unchanged pixels. These are metrics that properly quantify the trade-off between over and under detection.
Visual Evaluation
Using the images presented in Figures 6 and 7, quantitative change detection results were obtained and presented in Figures 9 and 10 for Beirut and Austin, respectively. Pixels in green correspond to true positives, in red to false positives, and in blue to false negatives. The first, second, and third rows individually contain change detection results generated using pre-event images: a Sentinel-II image (Figs. 9a and 10a), a synthetic Planet Dove image using the Regression + $\mathbf{W},\mathbf{C}$ model (Figs. 9b and 10b), described in the previous section, and a synthetic Planet Dove image generated with the proposed DDM-based algorithm (Figs. 9c and 10c). The Coloring procedure executed in the generation of the synthetic images made use of color information extracted from the post-Planet Dove image patches. A threshold $\omega$ has been chosen for each generated change map such that $\text{DR}=0.75$, which is an operating point that highlights the effects of the proposed algorithm.
(a) Sentinel-II (Pre) + Planet (Post)
(b) Regression + $\mathbf{W,C}$ Synthetic Planet (Pre) + Planet (Post)
(c) Proposed Synthetic Planet (Pre) + Planet (Post)
(d) Ground Truth—Planet Pre and Post
Figure 10: Change detection results in Austin Region for a fixed detection rate of 0.75. Pixels in green, blue, and red are true positives, false negatives, and false positives, respectively.
(a) Beirut Port Region
(b) Austin Region
Figure 11: Performance comparison of heterogeneous change detection region using a Planet Dove frame as a post-event image. Sentinel-II refers to when a Sentinel-II image is directly used as a pre-event image. Regression + W refers to when using the regression-based model with the whitening operation to generate a synthetic pre-event image. Synthetic Planet refers to when a synthetic Planet Dove image generated by the proposed I2I method is used as a pre-event image.
Comparing the results in Fig. 9, the proposed method has been able to heavily suppress false alarms. In fact, the change maps for Sentinel-II and Regression + W as pre-images show a staggering amount of false alarms, greatly surpassing the number of true positives. When the proposed method is applied, however, the majority of such false alarms are completely avoided. This visual superiority is further confirmed by the calculated metrics. For the Beirut region in Fig. 9, with $\text{DR}=0.75$, the proposed method reached a FAR of 0.03933 ($-log_{10}(\text{FAR})=-1.405$), whereas when using the Sentinel-II pre-event image, and the Regression + $\mathbf{W}$ synthetic pre-event image, the FAR was of 0.0731 ($-log_{10}(\text{FAR})=-1.136$) and 0.1055 ($-log_{10}(\text{FAR})=-0.9769$), respectively.
For the Austin region in Fig. 10, the improvements are less visually evident but still numerically expressive. In Fig. 10a, when the Sentinel-II image is used as pre-event image, multiple false alarms related to unchanged buildings can be spotted, specially in the bottom of the image, which are not observed in the change map for the proposed method’s change map (Fig. 10c). Moreover, for the regression-based change detection in Fig. 10b, a big false alarm is spotted in the eastern part of the image, which is not present in the change map of the proposed method. The metrics again prove the superiority of our technique: for $\text{DR}=0.75$, using the synthetic image generated by the proposed algorithm as the pre-event image leads to a FAR of 0.001678 ($-log_{10}(\text{FAR})=-2.775$), whereas when using Sentinel-II leads to a FAR of 0.007194 ($-log_{10}(\text{FAR})=-2.143$), and to a FAR of 0.003648 ($-log_{10}(\text{FAR})=-2.438$) when Regression + $\mathbf{W}$ is used.
Quantitative Evaluation
To better evaluate the performance of the proposed algorithm for all possible choices of clipping threshold $\omega$, employed in Line 10 of Algorithm 5, Figures 11a and 11b present receiving operating characteristic (ROC) comparisons between the baseline change detection and change detection using the proposed method, for Beirut port and Austin regions, respectively. These curves were obtained by varying $\omega$ between 0 and 1.0, with steps of 0.01. As can be visualized in the curves, $\omega$ adjusts how permissive or restrictive should the algorithm be towards potential changes: a trade-off between higher detection rates and lower false alarm rates.
In Fig. 11a, the proposed method beats both compared methods for all choices of $\omega$, which displays how it has been able to significantly reduce false alarms for a desired detection rate. For fixed detection rates, the difference between the false alarm rate of the proposed method and the other two cases is up to 0.7 on the logarithmic scale or 5 times lower on the linear scale (DR=0.45).
Meanwhile, in Fig. 11b, the performance of the proposed method is on par with the other two methods up to $-log_{10}(\text{FAR})=2$, displaying a slightly lower DR compared to the regression-based model: a maximum difference of 0.01 in the DR. After this point, however, the proposed method beats by a large margin the other two approaches, reaching a FAR difference of up to 1.25 in the logarithmic scale, or 17.8 times lower in the linear scale.
VI Discussion
The results of our study demonstrate the efficacy of leveraging Denoising Diffusion Implicit Models (DDIM) for large-scale optical image translation. Our approach successfully addresses the limitations of existing Image-to-Image (I2I) methods, particularly in handling large-scale multi-patch imagery. By focusing on super-resolving low spatial resolution images into high-resolution equivalents, our method achieves significant improvements in both radiometric accuracy and feature representation.
One of the key findings is the ability of our method to maintain uniformity across hundreds of patches, which is crucial for applications such as heterogeneous change detection (HCD). The comparative study against the standard classifier-free guided DDIM framework and other leading methods highlights the superior performance of our approach in terms of domain adaptation and artifact reduction. The HCD tasks in urban settings like Beirut, Lebanon, and Austin, USA, further validate the practical applicability of our method.
However, it is important to highlight that the simple change detection algorithm employed in our experiments does not fully exploit the enhanced visual resolution provided by our image translation method. This limitation is partly due to the algorithm’s design, which does not leverage the higher-resolution features effectively. Consequently, while our method shows promise, it does not necessarily outperform other methods, such as ShuffleMixer, SRDenseNet, and SwinIR, in HCD tasks using the chosen change detection algorithm. Despite their poor super-resolution performance, these methods maintain patch-wise feature consistency. The HCD experiments in this study are intended to demonstrate the potential utility of our approach rather than to establish it as the definitive solution for change detection. Future work will explore more advanced learning-based change detection procedures that can fully utilize the synthetic images generated by our method.
While our method effectively handles large-scale imagery, the computational complexity remains challenging. Future work will also explore optimization techniques to reduce the computational load without compromising the model’s performance. Additionally, extending our approach to other types of remote sensing data, such as radar or hyperspectral images, could broaden its applicability and impact. Another potential direction for future work is integrating our method with real-time processing systems. This would enable the rapid analysis of remote sensing data, particularly valuable in time-sensitive applications such as disaster response and environmental monitoring. Furthermore, incorporating advanced machine learning techniques, such as reinforcement learning, could enhance the adaptability and robustness of our method in diverse and dynamic environments.
In conclusion, this paper presented a novel and effective DDM-based approach for large-scale optical image translation addressing key challenges. The results demonstrate significant improvements in image quality and consistency, paving the way for more accurate and reliable remote sensing applications.
VII Conclusion
[…]
Appendix A Implementation Details
TABLE III: Hyperparameter values and their purpose for the performed experiments.
Parameter Value Purpose
Patch Size $128\times 128$ Width $w$ and height $h$ of the image patches fed to the neural network.
$N_{\text{final}}$ 64 Number of DDIM iterations for the final reverse diffusion in Line 16 of Alg. 3.
$N_{\text{pre}}$ 8 Number of DDIM iterations for the PSNR voting procedure in Line 15 of Alg. 3.
$n_{\text{noisy}}$ 8 Number of DDIM runs for the estimation of $\bm{\varepsilon}_{T}^{\text{best}}$, used in Line 13 of Alg. 3.
$\bm{\gamma}$ $\left[\cos\left(\frac{t_{i}/T+0.008}{1.008}\frac{\pi}{2}\right)^{2}\right]_{t_{i}}$ Diffusion noise level parameter for each timestep $t_{i}\in\{0,1,\dots,T-1\}$ [Nichol2021].
$T$ $1024$ Number of forward diffusion noising steps.
$p_{\text{uncond}}$ $10^{-1}$ Classifier-free guidance unconditional diffusion probability [ho2022classifierfree].
$\omega_{\text{uncond}}$ $1$ Classifier-free guidance inference weighting parameter [ho2022classifierfree].
$\lambda_{\text{consist}}$ $10^{-1}$ Weight of the consistency loss over the final loss [Daras2023].
$\varepsilon_{\text{consist}}$ $10$ It tells how far from the sampled timestep $t$ can the consistency timestep $t^{\prime}$ be sampled [Daras2023]. That is, $t^{\prime}\in\{t-\varepsilon_{\text{consist}},\dots,t\}$.
$n_{\text{consist}}$ $1$ Number of consistency training steps to be executed in a single overall training step [Daras2023].
$\gamma_{\text{SNR}}$ $5$ Clamping parameter for the Min-SNR-$\gamma$ training weighting strategy.
The model chosen for the experiments is the one presented in [Wang2022], where the authors propose a U-Net-style model that uses attention mechanisms to improve image fidelity. Differently from the diffusion-based models presented in [Saharia2021, Saharia2021SR3, ho2022classifierfree], this model processes the noisy input $\bm{y}_{t}$ and the conditioning input $\bm{x}$ independently. That is $\bm{y}_{t}$ is fed right into the encoder and is processed normally, whereas $\bm{x}$ is re-introduced into the decoder by inserting it into each residual block. To properly match the spatial resolution required by each residual block, $\bm{x}$ is processed by spatially adaptive normalization (SPADE) blocks. As confirmed by the performed experiments, we observed that this modification improved significantly the generation quality and fidelity to the conditioning input. Some internal parameters of the model were adjusted to maximize performance while limiting its size due to hardware constraints. The number of channels has been set to be a multiple of $96$. The number of residual blocks for each downsampling operation has been set to $2$. Attention resolutions were set to $2,4,8$ and $16$. Finally, the number of attention heads has been set to $4$. This configuration resulted in a model with around 195 million trainable parameters. For training, batch size has been set to $5$, and the learning rate has been set to $10^{-4}$.
We use the Huber loss function [meyer2020alternative] with its parameter $\delta$ set to $0.5$ since it has been shown to reduce the impact of outliers in the gradients while maintaining the benefits of the mean square distance for non-outliers. Implementation details for the Huber loss are presented in Appendix B-A3. RAdam optimizer [liu2021variance] has been chosen due to its proven training speed and its low sensitivity to the learning rate choice. The training ran for $31$ epochs, which was more than enough for the loss function to converge completely.
To stabilize training and avoid catastrophic forgetting of features, Exponential Moving Average (EMA) has been used in several DDM-based methods [Ho2020, Saharia2021, Wang2022, Dhariwal2021]. With the same goal, we instead use Stochastic Weight Averaging (SWA) [Izmailov2018], which consists of averaging weights after the training loss shows convergence behavior. Differently from EMA, SWA does not give higher importance to the weights of more recent epochs but instead considers all epochs after the beginning of weight averaging to have the same importance for overall learning. Since we have observed that the loss function converged after a number of epochs, it was reasonable to consider that newer weight updates would not bring an overall improvement of the model, thus applying SWA instead of EMA made more sense. The work presented in [garipov2018loss] suggests that averaging weights of different epochs can indeed result in higher performance than when only using the last weights of a training experiment. We observed that training stabilized after 10 epochs, so we started the SWA process at this moment. Also, SWA suggests increasing the learning rate during the epochs where weight averaging takes place. However, we noticed that increasing the learning rate over $10^{-4}$ made training unstable. For this reason, it is kept constant during the whole process.
For the performed experiments, multiple hyperparameters not directly related to the model setup have been chosen. They are presented in Table III, along with their purpose. Some choices were based on a grid search, and others were based on the articles that defined them in the first place. Moreover, two recently proposed training techniques for denoising diffusion models were included in our experiments: Consistency Diffusion Loss and Min-SNR-$\gamma$ Weighting. They are further discussed in Appendices B-D and B-E. The training and experiments were performed using one NVIDIA RTX A6000 GPU.
A-A Model Pruning
To elucidate the model’s performance in relation to its size, we have included Table IV, which presents the effects of various unstructured pruning rates. The experiments, executed with parameters $N_{\text{final}}=16$, $N_{\text{pre}}=2$, and $n_{\text{noisy}}=2$, demonstrate that a light pruning—specifically removing 1/8 of the weights by magnitude—significantly improved the model’s performance metrics on the test set used in the Experiments section. This enhancement suggests that such pruning can mitigate overfitting, streamlining the model to better generalize under different data regimes.
However, applying more aggressive pruning rates of 1/4 and 1/2 led to a considerable degradation in performance, indicating that such high levels of pruning do not preserve the model’s ability to perform effectively. This outcome suggests that there is a threshold beyond which pruning negatively impacts the model’s capacity to learn and generalize.
Regarding numerical efficiency, despite the model being converted into a sparse format in PyTorch post-pruning with the intention of exploiting sparsity for computational speedups, no such acceleration has been observed. The expected efficiency gains have not materialized, possibly due to the hardware’s inability to fully exploit the model’s sparsity properties. Another possibility is that the sparsity of the pruned models is not sufficiently large to result in noticeable speedups.
A-B Discussion on the Selection of Hyperparameters
In response to a possible question on the absence of ablation studies for the network parameters, it is important to note that the original paper presenting the network did not perform such studies either. The reason for this is twofold: firstly, diffusion models require extensive computational resources and time to train, often taking several days, even on high-end hardware. In our case, the hardware constraints were even more pronounced, which made the training process particularly challenging. Secondly, the potential gains from conducting ablation studies on different hyperparameters are not evidently clear. Diffusion models are complex, and their performance is not solely dependent on individual hyperparameters but rather on the intricate interplay between them. As such, varying single parameters may not result in significant changes in performance, making the exhaustive process of ablation less justifiable.
Furthermore, the chosen hyperparameters were set based on a combination of grid search results and recommendations from the literature, ensuring a well-informed starting point for our experiments. The internal parameters were carefully adjusted to balance performance optimization with the limitations imposed by our hardware. For instance, the number of channels, residual blocks, attention resolutions, and attention heads were all set to values that would maximize the model’s efficiency without exceeding our computational budget.
The training process itself was meticulously monitored, with the batch size and learning rate being set to values that allowed for stable and complete convergence of the loss function within 31 epochs. The use of the Huber loss function and the RAdam optimizer further contributed to the robustness and speed of the training, as detailed in Appendix B-A3.
In light of these considerations, we believe that the chosen approach for parameter selection and training was the most appropriate given the constraints and the nature of diffusion models. The inclusion of novel training techniques such as Consistency Diffusion Loss and Min-SNR-$\gamma$ Weighting further underscores our commitment to leveraging the latest advancements in the field to enhance our model’s performance.
TABLE IV: Model Performance Under Different Pruning Rates for $N_{\text{final}}=16$, $N_{\text{pre}}=2$, $n_{\text{noisy}}=2$
[…]
No Pruning 0.2229 52.46 14.60
1/8 0.2007 45.50 14.81
1/4 0.2636 66.33 13.85
1/2 0.4325 245.1 10.47
Appendix B Fundamental Concepts
B-A Conditional Denoising Diffusion Probabilistic Models
B-A1 Forward Diffusion
Let $\bm{y}_{0}$ be a data matrix sampled from $q(\bm{y}_{0})$. The forward diffusion process takes place by progressively adding Gaussian noise to $\bm{y}_{0}$. It is a Markov chain with $T$ transitions, or steps, described by the following conditional probability functions:
$q(\bm{y}_{1:T}|\bm{y}_{0})=\prod_{t=1}^{T}q(\bm{y}_{t}|\bm{y}_{t-1})$ (1)
$q(\bm{y}_{t}|\bm{y}_{t-1})=\mathcal{N}\left(\sqrt{\alpha_{t}}\bm{y}_{t-1},(1-\alpha_{t})\mathbf{I}\right)\text{.}$ (2)
It can be marginalized for each step as:
$q(\bm{y}_{t}|\bm{y}_{0}):=\int q(\bm{y}_{1:t}|\bm{y}_{0})\>\mathrm{\mathbf{d}}\bm{y}_{1:(t-1)}$ (3)
$=\mathcal{N}\left(\sqrt{\gamma_{t}}\bm{y}_{0},(1-\gamma_{t})\mathbf{I}\right)$ (4)
where $\gamma_{t}\in(0,1)$ is the noise schedule parameter at time $t$, and $\alpha_{t}=\gamma_{t}/\gamma_{t-1}$.
B-A2 Reverse Diffusion
Applying Bayes’, the posterior probability $q(\bm{y}_{t-1}|\bm{y}_{0},\bm{y}_{t})$ can be expressed as [luo_understanding_2022]:
$q(\bm{y}_{t-1}|\bm{y}_{0},\bm{y}_{t})=\mathcal{N}(\bm{\mu}_{q},\bm{\Sigma}_{q})$ (5)
$\bm{\mu}_{q}=\frac{1}{1-\gamma_{t}}\left[\sqrt{\gamma_{t-1}}(1-\alpha_{t})\bm{y}_{0}+\sqrt{\alpha_{t}}(1-\gamma_{t-1})\bm{y}_{t}\right]$ (6)
$\bm{\Sigma}_{q}=\dfrac{(1-\gamma_{t-1})(1-\alpha_{t})}{1-\gamma_{t}}\text{.}$ (7)
This means that it is possible to draw a sample from a previous diffusion step $t-1$ if $\bm{y}_{t}$ and $\bm{y}_{0}$ are known. Since the denoised data $\bm{y}_{0}$ is not known—as it is exactly what we want to retrieve at the end of the process—a neural network $\bm{\theta}_{t}$ can learn to predict an approximation of $\bm{y}_{0}$, $\hat{\bm{y}}_{0}$, given $\bm{y}_{t}$ and $\gamma_{t}$. In the context of conditional models, a conditional input $\bm{x}$ is also fed to the neural network.
From Eq. 4,
$\bm{y}_{t}=\sqrt{\gamma_{t}}\bm{y}_{0}+\sqrt{1-\gamma_{t}}\bm{\epsilon}_{t}$ (8)
$\text{where }\bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})\text{.}$ (9)
Thus, $\bm{y}_{0}$ can be parameterized as a linear combination between $\bm{y}_{t}$ and $\bm{\epsilon}_{t}$:
$\bm{y}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\bm{y}_{t}-\sqrt{1-\gamma_{t}}\bm{\epsilon}_{t}\right)\text{.}$ (10)
This parametrization is useful since Ho et al. [Ho2020] verified experimentally that higher quality approximations of $\bm{y}_{0}$ are obtained if the neural network is trained to instead predict $\bm{\epsilon}_{t}$, i.e., $\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})=\hat{\bm{\epsilon}}_{t}$. This prediction can be plugged into Eq. 10 to obtain $\hat{\bm{y}}_{0}$. Finally, with $\bm{y}_{t}$ and $\hat{\bm{y}}_{0}$, an approximation of $\bm{y}_{t-1}$ is obtained by sampling from Eq. 5. Combining Eqs. 5 and 10, such sampling procedure can be expressed as:
$\hat{\bm{y}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left[\bm{y}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\gamma_{t}}}\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})\right]$
$+\eta\sqrt{\dfrac{(1-\gamma_{t-1})(1-\alpha_{t})}{1-\gamma_{t}}}\bm{z}$ (11)
$\text{where }\bm{z}\sim\mathcal{N}(0,\mathbf{I})$ (12)
and $\text{and }\eta\in[0,1]\text{.}$ (13)
By repeating the reverse sampling until $t=1$, a higher quality $\hat{y}_{0}$ can be obtained.
B-A3 Training Objective
As discussed, the neural network should be trained to predict $\bm{\epsilon}_{t}$. Typically, the Mean Squared Error (MSE) serves as the default loss function [Ho2020]. However, it is highly sensitive to outliers, which can negatively affect the loss. To mitigate this, Huber Loss can be used, as it operates like MSE loss for regular data but switches to L1 loss for outliers.
Let $\delta\in\mathds{R}_{++}$ be the thresholding parameter. The training objective for $\bm{\theta}$ at timestep $t$ is defined as:
$\mathcal{L}_{\delta}(\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x}),\bm{\epsilon}_{t})=$
$\begin{cases}\frac{1}{2}(\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})-\bm{\epsilon}_{t})^{2}&\text{for }|\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})-\bm{\epsilon}_{t}|\leq\delta,\\ \delta(|\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})-\bm{\epsilon}_{t}|-\frac{1}{2}\delta)&\text{otherwise.}\end{cases}$ (14)
Finally, $\mathcal{L}_{\delta}$ is flattened and averaged to obtain a single loss value:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\ell_{\delta,i}$ (15)
where $\ell_{\delta,i}$ is the $i$-th element in the flattened $\mathcal{L}_{\delta}$ vector and $N$ is the number of elements in $\mathcal{L}_{\delta}$.
B-B Denoising Diffusion Implicit Models
Song et al. [Song2020] proposed Denoising Diffusion Implicit Models (DDIM). It consists of jumping steps in the reverse diffusion process to save computing resources while maintaining acceptable sampling quality. They observed that, when reducing the number of steps, setting $\eta=0$ in Eq. 13 yielded higher quality samples. This renders the reverse diffusion process deterministic, i.e., given fixed inputs, the output of the reverse diffusion is always the same. DDIM redefines the noise scheduling sequence $\bm{\gamma}$ by selecting a subset of the $T$ defined timesteps. In our experiments, we follow their linear selection procedure, which consists of selecting equally-spaced points of $\bm{\gamma}$. Such spacing between points is defined by the total number of desired sampling steps $T^{\text{DDIM}}$, which is a divisor of the original sampling steps $T$. Thus, the new noise schedule sequence $\bm{\gamma}^{\text{DDIM}}$ is defined as
$\gamma^{\text{DDIM}}_{t}=\gamma_{at}\>\>\>\forall\>0\leq t<T^{\text{DDIM}}$ (16)
where $\text{where }a=\frac{T^{\text{DDIM}}}{T}\text{.}$ (17)
B-C Classifier-Free Guidance
Classifier-free guidance is an approach that enhances the controllability of conditional diffusion models. Introduced in [ho2022classifierfree], it allows driving the reverse diffusion process towards desired data characteristics using input conditional data, as conditional diffusion models tend to ignore the conditioning input and produce samples that disregard the condition. The key idea is to train the model in a way that it can operate in two modes: one with guidance—conditioned on some input data—and one without—unconditioned. This is achieved by randomly dropping the conditioning input during training, which encourages the model to learn an internal representation that can be manipulated at sampling time to control the generation.
During training, $\bm{x}$ is dropped with probability $p$, effectively setting $\bm{x}=\O$. At sampling time, a guidance scale $\omega$ is introduced to amplify or diminish the influence of $\bm{x}$ by modifying the noise prediction $\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})=\hat{\bm{\epsilon}}_{t}$:
$\tilde{\bm{\epsilon}}_{t}=(1+\omega)\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})-\omega\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\O)\text{.}$ (18)
The modified noise matrix replaces $\bm{\theta}_{t}(\bm{y}_{t},\gamma_{t},\bm{x})$ in Eq. 13.
B-D Consistent Diffusion Models
The concept of Consistent Diffusion Models [Daras2023] addresses the challenge of sampling drift in denoising diffusion models. It focuses on enforcing a consistency property on the model during training. This approach comes from the observation that if $\bm{\epsilon}$ is fixed, the model should produce similar—consistent—predictions for similar diffusion timesteps. Following this line of thought, such consistency can be enforced by adding an additional loss term to the training loss.
Let $t$ and $t^{\prime}$ be diffusion timesteps such that $|t-t^{\prime}|<\varepsilon_{\text{consist}}$. The timestep $t^{\prime}$ is uniformly sampled from the interval $[t-\varepsilon_{\text{consist}},t]$. Following the line of thought of [Daras2023], the following loss term is summed to the diffusion loss:
$\mathcal{L}_{\text{consist}}=\lambda_{\text{consist}}\mathcal{L}\left(\hat{\bm{y}}_{0,t},\hat{\bm{y}}_{0,t^{\prime}}\right)$ (19)
where $\mathcal{L}$ is the chosen diffusion loss function, $\hat{\bm{y}}_{0,t}$ and $\hat{\bm{y}}_{0,t^{\prime}}$ are the estimations of $\bm{y}_{0}$ at diffusion steps $t$ and $t^{\prime}$ respectively. For a single training step, $n_{\text{consist}}$ timesteps are sampled from $[t-\varepsilon_{\text{consist}},t]$ and thus $n_{\text{consist}}$ consistent loss terms are added to the diffusion loss.
B-E Min-SNR Weighting Strategy
The Min-SNR Weighting Strategy [Hang2023] tackles the issue of slow training convergence in denoising diffusion models caused by conflicting optimization directions across different timesteps. To harmonize these trajectories and accelerate convergence, a clamped signal-to-noise ratio (SNR) is used to adaptively assign loss weights to each timestep. This method effectively suppresses the gradient conflicts among timesteps, leading to a more efficient training process.
Let $t$ be a diffusion timestep whose associated noise schedule value is $\gamma_{t}$, and $\gamma_{\text{SNR}}\in\mathds{R}_{++}$ be a clamping parameter. For the denoising diffusion objective, the Min-SNR weighting strategy defines the loss weight for timestep $t$ as:
$w_{t}=\min\left(\frac{\gamma_{\text{SNR}}}{\text{SNR}_{t}},1\right)$ (20)
$\text{where }\text{SNR}_{t}=\frac{\gamma_{t}}{1-\gamma_{t}}\text{.}$ (21)
This weighting scheme ensures that the model does not focus excessively on training samples with low noise values, which results in a significant convergence speedup, as observed in the experiments presented in [Hang2023]. The calculated weights are multiplied by loss values for each sample in a training batch.